OpenStudy (anonymous):

What is the solution set of |x - 6| = 10?

OpenStudy (anonymous):

solve two equations \[x-6=10\] and \[x-3=-10\]

OpenStudy (anonymous):

sorry second one is \[x-6=-10\]

OpenStudy (anonymous):

first one gives \[x=16\] second one gives \[x=-4\]

OpenStudy (anonymous):

Got it,thank you

OpenStudy (anonymous):

What is the solution set of |x| > 12?